OpenStudy (anonymous):

A basket contains the following pieces of fruit: 3 apples, 2 oranges, 2 bananas, 2 pears, and 5 peaches. Jonas picks a fruit at random and does not replace it. Then Beth picks a fruit at random. What is the probability that Jonas gets a peach and Beth gets an apple? a. 15/182 b. 8/27 c. 4/91 d. 15/196

jimthompson5910 (jim_thompson5910):

What is the probability that Jonas gets a peach?

OpenStudy (anonymous):

I don't know.. I don't get this. but I need it to be done.

jimthompson5910 (jim_thompson5910):

how many peaches are there?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

how many pieces of fruit are there

OpenStudy (anonymous):

14

jimthompson5910 (jim_thompson5910):

so the the probability that Jonas gets a peach is 5/14

jimthompson5910 (jim_thompson5910):

he takes the peach out, and does NOT replace it so there are 14-1 = 13 pieces of fruit left

OpenStudy (anonymous):

Ok and Beth is 3/13?

jimthompson5910 (jim_thompson5910):

there are still 3 apples, so the probability of Beth getting an apple is 3/13

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

multiply the two probabilities (5/14)*(3/13) = ???

OpenStudy (anonymous):

15/182

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

A true-false test has 12 questions. What is the probability of guessing the correct answer to all of the questions? a. 1/4096 b. 1/144 c. 1/24 d. 1/14 could you help with one more?

jimthompson5910 (jim_thompson5910):

There are 2 choices per question (T or F)

OpenStudy (anonymous):

does that mean you double 12?

jimthompson5910 (jim_thompson5910):

So there are 2^12 = 4,096 ways to answer the test (this is the number of combos of TFTFF...etc etc)

jimthompson5910 (jim_thompson5910):

There's only one way to get all 12 questions correct So that's why the probability is 1/4096

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

here's a smaller example say there are 3 questions. There are 3 ways to answer the test and they are TTT TTF TFT TFF FTT FTF FFT FFF So the probability of getting all 3 correct is 1/8

jimthompson5910 (jim_thompson5910):

notice how 2^3 = 8 that's not a coincidence

jimthompson5910 (jim_thompson5910):

oops meant to say "8 ways" not "3 ways"

OpenStudy (anonymous):

Ok, I'm starting to get it

jimthompson5910 (jim_thompson5910):

That's good